How much sodium chloride is needed to make 1 litre of an aqueous 2m solution? Mw = 58.5


1
Expert's answer
2021-02-15T03:19:29-0500

Q163450

How much sodium chloride is needed to make 1 liter of an aqueous 2M solution?

Molecular weight of NaCl = 58.5g/mol.


Solution:


We have to prepare 1Liter of 2M NaCl solution.


Volume of solution = 1 L.

Molarity of solution = 2 M.


Step 1: To find the moles of NaCl using the molarity formula.


Plug the volume and molarity in the formula of molarity and find the moles of NaCl required for preparing the given solution.


"Molarity = \frac{moles \space of \space NaCl }{volume \space of \space solution \space in \space 'L' }"



"2\space M = \frac{moles \space of \space NaCl }{1\space L}"


multiply both the side by 1L we have



"2\space mol/L * 1L = \frac{moles \space of \space NaCl }{1\space L } * 1L ; ........... (since \space M = mol/L)"


moles of NaCl = 2mol .


Step 2: Convert ' moles of NaCl' to grams by using the molar mass of NaCl.


The molar mass of NaCl = 58.5 g/mol.

This can be used as a conversion factor for inter- converting moles and grams.


1 mol NaCl = 58.5 g NaCl .



Grams of NaCl "= 2 \space mol \space NaCl * \frac{58.5 g \space of \space NaCl }{1 \space mol \space NaCl } = 117grams\space of\space NaCl"



Hence for preparing 1 L of 2 M NaCl solution, we will require 117 grams of NaCl.


I hope that this will help you understand the problem.

Please let me know if you have not understood any step.

Thank you. :)